\begin{thebibliography}{28}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Behjati et~al.(2019)Behjati, Moosavi-Dezfooli, Baghshah, and
  Frossard}]{behjati2019universal}
Melika Behjati, Seyed-Mohsen Moosavi-Dezfooli, Mahdieh~Soleymani Baghshah, and
  Pascal Frossard. 2019.
\newblock Universal adversarial attacks on text classifiers.
\newblock In \emph{ICASSP 2019-2019 IEEE International Conference on Acoustics,
  Speech and Signal Processing (ICASSP)}, pages 7345--7349. IEEE.

\bibitem[{Carlini and Wagner(2017)}]{carlini2017towards}
Nicholas Carlini and David Wagner. 2017.
\newblock Towards evaluating the robustness of neural networks.
\newblock In \emph{2017 IEEE Symposium on Security and Privacy (SP)}, pages
  39--57. IEEE.

\bibitem[{Cisse et~al.(2017)Cisse, Adi, Neverova, and
  Keshet}]{cisse2017houdini}
Moustapha Cisse, Yossi Adi, Natalia Neverova, and Joseph Keshet. 2017.
\newblock Houdini: Fooling deep structured prediction models.
\newblock \emph{arXiv preprint arXiv:1707.05373}.

\bibitem[{Devlin et~al.(2018)Devlin, Chang, Lee, and
  Toutanova}]{devlin2018bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018.
\newblock Bert: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock \emph{arXiv preprint arXiv:1810.04805}.

\bibitem[{Ebrahimi et~al.(2017)Ebrahimi, Rao, Lowd, and
  Dou}]{ebrahimi2017hotflip}
Javid Ebrahimi, Anyi Rao, Daniel Lowd, and Dejing Dou. 2017.
\newblock Hotflip: White-box adversarial examples for text classification.
\newblock \emph{arXiv preprint arXiv:1712.06751}.

\bibitem[{Goodfellow et~al.(2014)Goodfellow, Shlens, and
  Szegedy}]{goodfellow2014explaining}
Ian~J Goodfellow, Jonathon Shlens, and Christian Szegedy. 2014.
\newblock Explaining and harnessing adversarial examples.
\newblock \emph{arXiv preprint arXiv:1412.6572}.

\bibitem[{Grosse et~al.(2016)Grosse, Papernot, Manoharan, Backes, and
  McDaniel}]{GrossePM0M16}
Kathrin Grosse, Nicolas Papernot, Praveen Manoharan, Michael Backes, and
  Patrick~D. McDaniel. 2016.
\newblock \href {http://arxiv.org/abs/1606.04435} {Adversarial perturbations
  against deep neural networks for malware classification}.
\newblock \emph{CoRR}, abs/1606.04435.

\bibitem[{Jia and Liang(2017)}]{jia2017adversarial}
Robin Jia and Percy Liang. 2017.
\newblock Adversarial examples for evaluating reading comprehension systems.
\newblock \emph{arXiv preprint arXiv:1707.07328}.

\bibitem[{Kingma and Ba(2014)}]{kingma2014adam}
Diederik~P Kingma and Jimmy Ba. 2014.
\newblock Adam: A method for stochastic optimization.
\newblock \emph{arXiv preprint arXiv:1412.6980}.

\bibitem[{Kurakin et~al.(2016{\natexlab{a}})Kurakin, Goodfellow, and
  Bengio}]{kurakin2016adversarial_e}
Alexey Kurakin, Ian Goodfellow, and Samy Bengio. 2016{\natexlab{a}}.
\newblock Adversarial examples in the physical world.
\newblock \emph{arXiv preprint arXiv:1607.02533}.

\bibitem[{Kurakin et~al.(2016{\natexlab{b}})Kurakin, Goodfellow, and
  Bengio}]{kurakin2016adversarial_a}
Alexey Kurakin, Ian Goodfellow, and Samy Bengio. 2016{\natexlab{b}}.
\newblock Adversarial machine learning at scale.
\newblock \emph{arXiv preprint arXiv:1611.01236}.

\bibitem[{Li et~al.(2018)Li, Ji, Du, Li, and Wang}]{li2018textbugger}
Jinfeng Li, Shouling Ji, Tianyu Du, Bo~Li, and Ting Wang. 2018.
\newblock Textbugger: Generating adversarial text against real-world
  applications.
\newblock \emph{arXiv preprint arXiv:1812.05271}.

\bibitem[{Liang et~al.(2017)Liang, Li, Su, Bian, Li, and Shi}]{liang2017deep}
Bin Liang, Hongcheng Li, Miaoqiang Su, Pan Bian, Xirong Li, and Wenchang Shi.
  2017.
\newblock Deep text classification can be fooled.
\newblock \emph{arXiv preprint arXiv:1704.08006}.

\bibitem[{Moosavi-Dezfooli et~al.(2017{\natexlab{a}})Moosavi-Dezfooli, Fawzi,
  Fawzi, and Frossard}]{moosavi2017universal}
Seyed-Mohsen Moosavi-Dezfooli, Alhussein Fawzi, Omar Fawzi, and Pascal
  Frossard. 2017{\natexlab{a}}.
\newblock Universal adversarial perturbations.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 1765--1773.

\bibitem[{Moosavi-Dezfooli et~al.(2017{\natexlab{b}})Moosavi-Dezfooli, Fawzi,
  Fawzi, Frossard, and Soatto}]{moosavi2017analysis}
Seyed-Mohsen Moosavi-Dezfooli, Alhussein Fawzi, Omar Fawzi, Pascal Frossard,
  and Stefano Soatto. 2017{\natexlab{b}}.
\newblock Analysis of universal adversarial perturbations.
\newblock \emph{arXiv preprint arXiv:1705.09554}.

\bibitem[{Moosavi-Dezfooli et~al.(2016)Moosavi-Dezfooli, Fawzi, and
  Frossard}]{moosavi2016deepfool}
Seyed-Mohsen Moosavi-Dezfooli, Alhussein Fawzi, and Pascal Frossard. 2016.
\newblock Deepfool: a simple and accurate method to fool deep neural networks.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 2574--2582.

\bibitem[{Mopuri et~al.(2017)Mopuri, Garg, and Babu}]{mopurigb17}
Konda~Reddy Mopuri, Utsav Garg, and R.~Venkatesh Babu. 2017.
\newblock \href {http://arxiv.org/abs/1707.05572} {Fast feature fool: {A} data
  independent approach to universal adversarial perturbations}.
\newblock \emph{CoRR}, abs/1707.05572.

\bibitem[{Papernot et~al.(2016)Papernot, McDaniel, Swami, and
  Harang}]{papernot2016crafting}
Nicolas Papernot, Patrick McDaniel, Ananthram Swami, and Richard Harang. 2016.
\newblock Crafting adversarial input sequences for recurrent neural networks.
\newblock In \emph{MILCOM 2016-2016 IEEE Military Communications Conference},
  pages 49--54. IEEE.

\bibitem[{Papernot et~al.(2015)Papernot, McDaniel, Jha, Fredrikson, Celik, and
  Swami}]{PapernotMJFCS15}
Nicolas Papernot, Patrick~D. McDaniel, Somesh Jha, Matt Fredrikson, Z.~Berkay
  Celik, and Ananthram Swami. 2015.
\newblock \href {http://arxiv.org/abs/1511.07528} {The limitations of deep
  learning in adversarial settings}.
\newblock \emph{CoRR}, abs/1511.07528.

\bibitem[{Poursaeed et~al.(2017)Poursaeed, Katsman, Gao, and
  Belongie}]{omid_generative_2017}
Omid Poursaeed, Isay Katsman, Bicheng Gao, and Serge~J. Belongie. 2017.
\newblock \href {http://arxiv.org/abs/1712.02328} {Generative adversarial
  perturbations}.
\newblock \emph{CoRR}, abs/1712.02328.

\bibitem[{Ribeiro et~al.(2018)Ribeiro, Singh, and
  Guestrin}]{ribeiro2018semantically}
Marco~Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. 2018.
\newblock Semantically equivalent adversarial rules for debugging nlp models.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 856--865.

\bibitem[{Samanta and Mehta(2018)}]{samanta2018generating}
Suranjana Samanta and Sameep Mehta. 2018.
\newblock Generating adversarial text samples.
\newblock In \emph{Advances in Information Retrieval}, pages 744--749, Cham.
  Springer International Publishing.

\bibitem[{Shafahi et~al.(2018)Shafahi, Najibi, Xu, Dickerson, Davis, and
  Goldstein}]{shafahi2018universal}
Ali Shafahi, Mahyar Najibi, Zheng Xu, John Dickerson, Larry~S Davis, and Tom
  Goldstein. 2018.
\newblock Universal adversarial training.
\newblock \emph{arXiv preprint arXiv:1811.11304}.

\bibitem[{Sun et~al.(2018)Sun, Tang, Yi, Wang, and
  Zhou}]{Identify_Mengying_2018}
Mengying Sun, Fengyi Tang, Jinfeng Yi, Fei Wang, and Jiayu Zhou. 2018.
\newblock \href {http://arxiv.org/abs/1802.04822} {Identify susceptible
  locations in medical records via adversarial attacks on deep predictive
  models}.
\newblock \emph{CoRR}, abs/1802.04822.

\bibitem[{Szegedy et~al.(2016)Szegedy, Vanhoucke, Ioffe, Shlens, and
  Wojna}]{szegedy2016rethinking}
Christian Szegedy, Vincent Vanhoucke, Sergey Ioffe, Jon Shlens, and Zbigniew
  Wojna. 2016.
\newblock Rethinking the inception architecture for computer vision.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 2818--2826.

\bibitem[{Szegedy et~al.(2013)Szegedy, Zaremba, Sutskever, Bruna, Erhan,
  Goodfellow, and Fergus}]{szegedy2013intriguing}
Christian Szegedy, Wojciech Zaremba, Ilya Sutskever, Joan Bruna, Dumitru Erhan,
  Ian Goodfellow, and Rob Fergus. 2013.
\newblock Intriguing properties of neural networks.
\newblock \emph{arXiv preprint arXiv:1312.6199}.

\bibitem[{Wang et~al.(2018)Wang, Singh, Michael, Hill, Levy, and
  Bowman}]{wang2018glue}
Alex Wang, Amapreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel~R
  Bowman. 2018.
\newblock Glue: A multi-task benchmark and analysis platform for natural
  language understanding.
\newblock \emph{arXiv preprint arXiv:1804.07461}.

\bibitem[{Zhang et~al.(2019)Zhang, Sheng, and Alhazmi}]{Zhang2019GeneratingTA}
Wei~Emma Zhang, Quan~Z. Sheng, and Ahoud Abdulrahmn~F. Alhazmi. 2019.
\newblock \href {http://arxiv.org/abs/1901.06796} {Generating textual
  adversarial examples for deep learning models: {A} survey}.
\newblock \emph{CoRR}, abs/1901.06796.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{emnlp-ijcnlp-2019}
\usepackage{times}
\usepackage{latexsym}

\usepackage{url}
\usepackage{algorithm2e}
\aclfinalcopy 

\title{Universal Adversarial Perturbation for Text Classification}

\author{Hang Gao \\
  University of Maryland, \\ Baltimore County \\
  1000 Hilltop Cir, Baltimore, MD 21250 \\ 
  {\tt hanggao1@umbc.edu} \\\And
  Tim Oates \\
  University of Maryland, \\  Baltimore County \\
  1000 Hilltop Cir, Baltimore, MD 21250 \\
  {\tt oates@cs.umbc.edu} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
Given a state-of-the-art deep neural network text classifier, we show the existence of a universal and very small perturbation vector (in the embedding space) that causes natural text to be misclassified with high probability. Unlike images on which a single fixed-size adversarial perturbation can be found, text is of variable
length, so we define the ``universality'' as ``token-agnostic'', where a single perturbation is applied to each token, resulting in different perturbations of flexible sizes at the sequence level. We propose an algorithm to compute universal adversarial perturbations, and show that the state-of-the-art deep neural networks are highly vulnerable to them, even though they keep the neighborhood of
tokens mostly preserved. We also show how to use these adversarial perturbations to generate adversarial text samples. The surprising existence of universal ``token-agnostic'' adversarial perturbations may reveal important properties of a text classifier.

\end{abstract}

\section{Introduction}
Deep neural networks (DNNs) are vulnerable to adversarial samples, i.e., carefully crafted samples with imperceptible perturbations designed to mislead a pre-trained model \cite{szegedy2013intriguing}, raising security concerns.

Two types of adversarial perturbations can be considered: universal or sample-dependent. Sample-dependent perturbations can vary for different samples in a dataset, while universal ones remain fixed for all samples. Adversarial samples were first studied for image DNNs with various methods proposed to generate and defend both sample-dependent perturbations \cite{szegedy2016rethinking, carlini2017towards, goodfellow2014explaining, moosavi2016deepfool, kurakin2016adversarial_e, kurakin2016adversarial_a, cisse2017houdini} and universal perturbations \cite{moosavi2017universal, mopurigb17, moosavi2017analysis}. Later, research on attacking DNNs for text applications emerged \cite{liang2017deep, ebrahimi2017hotflip, samanta2018generating, jia2017adversarial, li2018textbugger, ribeiro2018semantically}, where the definition of an ``imperceptible perturbation'' can differ from image-based perturbations. 

Although most existing adversarial attack methods on textual deep learning models are sample-dependent, recently, \newcite{behjati2019universal} proposed universal adversarial attacks on text classifiers by generating a universal sequence of words that can be added to any input to fool a classifier with high probability. Their method is effective, but is unlikely to preserve the syntax or semantic of the original input text. In this paper, we instead seek to generate universal adversarial perturbations and samples for text classifiers, by explicitly limiting the norm of changes. 

Our main contributions can be summarized as follows:

\begin{itemize}
    \item We show that a textual deep model can be vulnerable to some small token-agnostic perturbation in the token embedding space.
    \item We propose an algorithm to compute such universal adversarial perturbations. 
    \item We provide a way to generate adversarial samples in textual form, given the universal perturbations. 
\end{itemize}

\section{Related Work}
In general, attack methods on textual deep learning models can be categorized based on the following five criteria \cite{Zhang2019GeneratingTA}: (1) model access; (2) application; (3) target; (4) granularity and (5) the attacked DNNs.

\subsection{White-Box Attack}
Considering model access, white-box attacks require access to a model's full information. 

FGSM \cite{goodfellow2014explaining} was the first attack on image-based models. TextFool \cite{liang2017deep} designed three perturbation strategies: insertion, modification, and removal to generate adversarial samples, based on the concept of FGSM. \newcite{samanta2018generating}, adopting the same idea as TextFool, proposed to craft adversarial text samples by deleting or replacing the important or salient words in the text or by introducing new words to the text sample. \newcite{papernot2016crafting} used the forward derivative such as JSMA \cite{PapernotMJFCS15} to find the sequence that contributes the most towards the adversary direction. \newcite{GrossePM0M16}, on the other hand, crafted adversarial samples on input features by computing the gradient of the model Jacobian to estimate the perturbation direction. \newcite{Identify_Mengying_2018} adopted the C\&W method \cite{carlini2017towards} for attacking predictive models of medical records. Other kinds of methods are summarized in \cite{Zhang2019GeneratingTA}.

\subsection{Universal Adversarial Perturbation}
Universal adversarial perturbation was first studied for image-based deep models by \cite{moosavi2017universal}. They showed the existence of universal image-agnostic perturbations with remarkable generalization properties. Later on, \newcite{omid_generative_2017} proposed to generate universal perturbations with neural networks. \newcite{shafahi2018universal}, on the other hand, introduced an algorithm for universal adversarial training. For natural language processing, \newcite{behjati2019universal} proposed to generate universal adversarial samples by finding a universal sequence of words that can be added to any input to fool a classifier.

\section{Proposed Methods}
Given a set of samples $\mathbf{X} = \{\mathbf{x}_i, i=1,...,N\}$, and a network $f(w, \cdot)$ with frozen parameter $w$ that maps each sequence of text $\mathbf{x}_i$ onto labels, we define a universal adversarial perturbation $\delta$ to optimize the following objective:

\begin{equation}
    \max\limits_{\delta}\; \frac{1}{N}\sum_{i=1}^N l(w, \mathbf{x}_i+\delta) \;\;\mathrm{s.t.}\;\; \|\delta\|_p \le \epsilon
\end{equation}


\noindent where $l(w, \cdot)$ represents the loss used for training DNNs. We use cross-entropy as the training loss in this paper. $\epsilon$ denotes the maximum $p$-norm of $\delta$.

\begin{table*}[h!]
\centering
 \begin{tabular}{|c c c c c c c|} 
 \hline
 $\mathrm{Corpus}$ & $|\mathrm{Train}|$ & $|\mathrm{Dev}|$ & $|\mathrm{Test}|$ & $\mathrm{Task}$ & $\mathrm{Metrics}$ & $\mathrm{Domain}$\\ 
 \hline
 MRPC & 3.7k & 0.4k & 1.7k & paraphrase & acc./F1 & news \\
 \hline 
 CoLA & 8.5k & 1k  & 1k & acceptability & Matthews corr. & misc. \\ 
 \hline 
 SST-2 & 67k & 0.8k & 1.8k & sentiment & acc. & movie reviews \\ 
 \hline 
 QNLI & 105k & 5.5k & 5.4k & QA/NLI & acc. & Wikipedia \\
 \hline 
 \end{tabular}
  \caption{The statistics of the datasets used for the experiments.}
\end{table*}

\begin{algorithm}[t]
\SetKwInOut{Input}{Input}
\SetKwInOut{Output}{Output}
\SetAlgoLined
\Input{ Training samples X, perturbation maximum norm $\epsilon$, learning rate $\lambda$}
\Output{ Universal adversarial perturbation $\delta$ }
 Initialize $\delta \leftarrow 0$\;
 \For{ $\mathrm{epoch} = 1,..., N_{ep}$}{
    \For{$\mathrm{minibatch }\; B \subset X$}{
        \eIf{$\mathrm{normalize\; gradients}$}{
            $g = \nabla_{\delta}l(w, x + \delta)$ \;
            $ng = -\epsilon \frac{g}{\|g\|_p}$ \;
            $\delta \leftarrow \delta + \lambda ng$ \;
        }{
            $g = \nabla_{\delta}l(w, x + \delta)$ \;
            $\delta \leftarrow \delta + \lambda g$ \;
        }
        $\mathrm{Project} \; \delta \; \mathrm{to} \; l_p \; \mathrm{ball}$ \;
    }
 }
 \caption{The proposed algorithm to generate a universal perturbation. }
\end{algorithm}

Given a sequence of text sample $\mathbf{x}_i = x_{i, 1}, ..., x_{i, m}$ of length $m$, with $e_{i, 1}, ..., e_{i, m}$ as the embeddings for each token, the corresponding perturbed sample $\mathbf{x}_i + \delta$ is generated by adding $\delta$ to each embedding, i.e. $e_{i, 1}+\delta, ..., e_{i, m}+\delta$. The perturbation is ``universal" as ``token-agnostic", i.e. it remains the same regardless of the token to which it is applied. We present the algorithm to generate universal adversarial perturbations in Algorithm 1. 

While our definition of adversarial perturbations is similar to the one for images, the ``imperceptibleness" is different as token embeddings, unlike images, usually have no direct semantic meaning for human beings. In this paper, we define ``imperceptibleness" as the normalized intersection of the neighborhoods of a token $t$ before and after perturbation ($N^{(t)}_b$ and $N^{(t)}_a$), i.e.

\begin{equation}
    NI(t) = \frac{|N^{(t)}_b \cap N^{(t)}_a|}{\min (|N^{(t)}_a|, |N^{(t)}_b|)}
\end{equation}

\noindent and the ``imperceptibleness" score for a vocabulary $V$ is then defined as,

\begin{equation}
    NI(V) = \frac{1}{|V|} \sum_{t \in V} NI(t)
\end{equation}

\section{Experiment}
\subsection{Baseline}
Although we are interested in a single universal perturbation applicable to all tokens, we establish a strong baseline where a unique adversarial perturbation is obtained for each token in a vocabulary. The ``universality" is then considered as ``sample-agnostic'' as perturbations are now token-sensitive, but remain unchanged for the same token in different samples. These perturbations can be generated in a similar way as the one presented in Algorithm 1.

We build another set of baselines by randomizing all perturbations, and then normalizing them to the maximum $p$-norm limit $\epsilon$. 

\subsection{Classifier and Dataset}
In this paper, we focus on generating universal adversarial perturbations for text classifiers. We adopt BERT \cite{devlin2018bert} as the classifier model for its promising performance on the task of text classification. We fine tune the BERT base model with its uncased tokenizer on four GLUE datasets \cite{wang2018glue}: MRPC, CoLA, SST-2 and QNLI. The details of these datasets are shown in Table 1. For each task, we explore hyper-parameters with the following ranges, and pick the classifier with the best performance on the evaluation data. 

\begin{itemize}
    \item \textbf{Batch size:} 16, 32
    \item \textbf{Learning rate:} 5e-5, 3e-5 and 2e-5
    \item \textbf{Number of epochs:} 3, 4
\end{itemize}

We then split the training data of each task into train/dev sets and train our universal adversarial perturbation and samples on the train set with hyper-parameters picked on the dev set. We test the performance of these perturbation and samples on the original evaluation data of each task, which is never used for parameter-tuning or hyper-parameter selection for those perturbation and samples. 

\subsection{Result}

\begin{table}[ht!]
\centering
 \begin{tabular}{|c c c c c|} 
 \hline
 \multicolumn{1}{|c}{Task} & \multicolumn{4}{c|}{Maximum p-Norm $\epsilon$} \\
 \hline
 MRPC$\mathrm{(0.882)}$ & $\mathrm{0.05}$  & $\mathrm{0.1}$ & $\mathrm{0.15}$ & $\mathrm{0.2}$\\ 
 \hline
 $\mathrm{Baseline}_{\mathrm{VR}}$ & 0.882 & 0.876 & 0.876 & 0.876  \\
 $\mathrm{Baseline}_{\mathrm{SR}}$ & 0.884 & 0.884 & 0.882 & 0.884  \\
 $\mathrm{Baseline}_{\mathrm{V}}$ & 0.863 & 0.858 & 0.831 & 0.820  \\
 $\mathrm{Ours}$ & \textbf{0.860} & \textbf{0.811} & \textbf{0.770} & \textbf{0.748}  \\
 \hline 
 CoLA$\mathrm{(0.585)}$ & $\mathrm{0.05}$  & $\mathrm{0.1}$ & $\mathrm{0.15}$ & $\mathrm{0.2}$\\ 
 \hline 
 $\mathrm{Baseline}_{\mathrm{VR}}$ & 0.593 & 0.591 & 0.588 & 0.588  \\
 $\mathrm{Baseline}_{\mathrm{SR}}$ & 0.586 & 0.588 & 0.583 & 0.578  \\
 $\mathrm{Baseline}_{\mathrm{V}}$ & 0.580 & 0.575 & 0.557 & 0.346  \\
 $\mathrm{Ours}$ & \textbf{0.580} & \textbf{0.566} & \textbf{0.278} & \textbf{0.168}  \\
 \hline 
 SST-2$\mathrm{(0.929)}$ & $\mathrm{0.05}$  & $\mathrm{0.1}$ & $\mathrm{0.15}$ & $\mathrm{0.2}$\\ 
 \hline 
 $\mathrm{Baseline}_{\mathrm{VR}}$ & 0.928 & 0.925 & 0.924 & 0.924  \\
 $\mathrm{Baseline}_{\mathrm{SR}}$ & 0.928 & 0.929 & 0.930 & 0.931  \\
 $\mathrm{Baseline}_{\mathrm{V}}$ &  0.925 & 0.876 & 0.750 & 0.608 \\
 $\mathrm{Ours}$ & \textbf{0.892} & \textbf{0.823} & \textbf{0.704} & \textbf{0.585}  \\
 \hline 
 QNLI$\mathrm{(0.918)}$ & $\mathrm{0.05}$  & $\mathrm{0.1}$ & $\mathrm{0.15}$ & $\mathrm{0.2}$\\ 
 \hline 
 $\mathrm{Baseline}_{\mathrm{VR}}$ & 0.918 & 0.917 & 0.917 & 0.917  \\
 $\mathrm{Baseline}_{\mathrm{SR}}$ & 0.917 & 0.917 & 0.916 & 0.915\\
 $\mathrm{Baseline}_{\mathrm{V}}$ & 0.916 & 0.917 & 0.915 & 0.908  \\
 $\mathrm{Ours}$ & \textbf{0.916} & \textbf{0.897} & \textbf{0.844} & \textbf{0.499}  \\
 \hline 
 \end{tabular}
  \caption{The comparison of the proposed methods with baseline approaches on the evaluation data for each task. The performance of the classifier without adversarial perturbation is presented in the parentheses. $\mathrm{Baseline}_{\mathrm{VR}}$: baseline with vocabulary based random perturbations; $\mathrm{Baseline}_{\mathrm{SR}}$: baseline with a single random perturbation; $\mathrm{Baseline}_{\mathrm{V}}$: baseline with vocabulary based adversarial perturbations.}
\end{table}

\begin{table}[t]
\centering
 \begin{tabular}{|c c c c c|} 
 \hline
 \multicolumn{1}{|c}{Task} & \multicolumn{4}{c|}{Maximum p-Norm $\epsilon$} \\
 \hline
 MRPC & $\mathrm{0.05}$  & $\mathrm{0.1}$ & $\mathrm{0.15}$ & $\mathrm{0.2}$\\ 
 \hline
 $\mathrm{Baseline}_{\mathrm{VR}}$ & 0.907 & 0.830 & 0.762 & 0.697  \\
 $\mathrm{Baseline}_{\mathrm{SR}}$ & \textbf{0.974} & \textbf{0.949} & \textbf{0.919} & \textbf{0.885} \\
 $\mathrm{Baseline}_{\mathrm{V}}$ & 0.897 & 0.811 & 0.731 & 0.658  \\
 $\mathrm{Ours}$ & 0.968 & 0.935 & 0.904 & 0.863  \\
 \hline 
 CoLA & $\mathrm{0.05}$  & $\mathrm{0.1}$ & $\mathrm{0.15}$ & $\mathrm{0.2}$\\ 
 \hline 
 $\mathrm{Baseline}_{\mathrm{VR}}$ & 0.920 & 0.853 & 0.789 & 0.726  \\
 $\mathrm{Baseline}_{\mathrm{SR}}$ & \textbf{0.972} & \textbf{0.944} & \textbf{0.913} & \textbf{0.879}  \\
 $\mathrm{Baseline}_{\mathrm{V}}$ & 0.906 & 0.815 & 0.740 & 0.690  \\
 $\mathrm{Ours}$ & 0.963 & 0.931 & 0.905 & 0.863  \\
 \hline 
 SST-2 & $\mathrm{0.05}$  & $\mathrm{0.1}$ & $\mathrm{0.15}$ & $\mathrm{0.2}$\\ 
 \hline 
 $\mathrm{Baseline}_{\mathrm{VR}}$ & 0.916 & 0.840 & 0.776 & 0.719  \\
 $\mathrm{Baseline}_{\mathrm{SR}}$ & \textbf{0.975} & \textbf{0.949} & \textbf{0.920} & \textbf{0.888}  \\
 $\mathrm{Baseline}_{\mathrm{V}}$ &  0.895 & 0.804 & 0.717 & 0.654 \\
 $\mathrm{Ours}$ & 0.954 & 0.913 & 0.871 & 0.835  \\
 \hline 
 QNLI & $\mathrm{0.05}$  & $\mathrm{0.1}$ & $\mathrm{0.15}$ & $\mathrm{0.2}$\\ 
 \hline 
 $\mathrm{Baseline}_{\mathrm{VR}}$ & 0.886 & 0.793 & 0.716 & 0.649  \\
 $\mathrm{Baseline}_{\mathrm{SR}}$ & \textbf{0.975} & \textbf{0.950} & \textbf{0.922} & \textbf{0.891}\\
 $\mathrm{Baseline}_{\mathrm{V}}$ & 0.866 & 0.761 & 0.678 & 0.608 \\
 $\mathrm{Ours}$ & 0.959 & 0.943 & 0.916 & 0.877  \\
 \hline 
 \end{tabular}
  \caption{The $NI(V)$ score on each task. The number of neighbors for each token is 5 and we use cosine similarity as the distance measure between any two tokens. }
\end{table}

We use Adam \cite{kingma2014adam} as the optimization algorithm and normalized gradient for the perturbation computation. For each task, we split the training data into train/eval sets with a 0.9/0.1 ratio, train the perturbation on the train set and early stop on the eval set. The train batch size is set to 32 and the eval batch size is set to 8. $p$ is set to 2 and the learning rate $\lambda$ is set to 0.05. 

We range the maximum $p$-norm of perturbation $\epsilon$ from 0.05 to 0.2 and present the perturbed classifier's performance on the evaluation data of each task in Table 2. The choice of $\epsilon$ is determined so that the norm of the perturbation is significantly smaller than the average norm of token embeddings. For each task, we compare the proposed approach with three different baselines: $\mathrm{Baseline}_{\mathrm{VR}}$ - vocabulary based perturbations that are randomly initialized with $p$-norm scaled to $\epsilon$; $\mathrm{Baseline}_{\mathrm{SR}}$ - a single perturbation that is randomly initialized with $p$-norm scaled to $\epsilon$ and $\mathrm{Baseline}_{\mathrm{V}}$ - vocabulary based adversarial perturbations trained similarly as Algorithm 1.

The results in Table 2 indicate that (1) randomly initialized perturbations can hardly fool the classifier in spite of potentially larger $p$-norm than the trained ones; (2) our proposed single universal adversarial perturbation outperforms all the baselines on all the tasks. 

To evaluate the ``imperceptibleness" of the perturbation, we show the ``imperceptibleness" score defined in Equation (3) in Table 3. The number of neighbors for each token is set to 5 and we use cosine as the distance measure between any two tokens. Single perturbation based methods are better at preserving the neighborhood of each token than vocabulary based ones. And our proposed method is comparable with $\mathrm{Baseline}_{\mathrm{SR}}$ with slightly lower performance.

\begin{table}[t]
\centering
 \begin{tabular}{|c c c c c|} 
 \hline
 \multicolumn{5}{|c|}{Data Ratio} \\
 \hline
 $\mathrm{10\%}$ & $\mathrm{20\%}$  & $\mathrm{30\%}$ & $\mathrm{40\%}$ & $\mathrm{50\%}$\\ 
 \hline
 0.663 & 0.592 & 0.663 & 0.592 & 0.661   \\
 \hline 
 $\mathrm{60\%}$ & $\mathrm{70\%}$  & $\mathrm{80\%}$ & $\mathrm{90\%}$ & \\ 
 \hline
 0.588 & 0.589 & 0.663 & 0.662 & \\ 
 \hline
 \end{tabular}
  \caption{The evaluation result (Accuracy) on SST-2 with the universal adversarial perturbation trained on different ratios of training data. 10\% of training data is enough to generate perturbation with good performance.}
\end{table}

\begin{table}[b]
\centering
 \begin{tabular}{|p{0.85\linewidth}|c|} 
 \hline 
 Text (top: original, bottom: adversarial) & Prediction \\
 \hline
 \textbf{the} bears sniffed & 1 \\ 
 \textbf{a} bears sniffed & 0 \\
 \hline 
 i walk \textbf{and} dana runs . & 1 \\ 
 i walk \textbf{,} dana runs . & 0 \\
 \hline
  sue gave to bill \textbf{a} book . & 1\\ 
  sue gave to bill \textbf{the} book . & 0 \\
 \hline
  terry delighted \textbf{.} & 0 \\
  terry delighted \textbf{;} & 1 \\
 \hline
 dana walking \textbf{and} leslie running . & 0 \\
 dana walking \textbf{,} leslie running . & 1 \\ 
 \hline
 \end{tabular}
  \caption{A few adversarial samples generated on CoLA evaluation data. $\epsilon$ is set to 0.2.}
\end{table}

We further evaluate how much data is required to generate the universal adversarial perturbation and present the result in Table 4. The performance is reported as Accuracy on the SST-2 dataset with various ratios (from 10\% to 90\%) of training data used to generate the perturbation. The maximum $p$-norm is set to 0.15 and we adopt SGD as the optimization algorithm with momentum set to 0.9 and gradients unnormalized. The result indicates that 10\% of training data is enough to generate perturbations with comparable performance, suggesting the scalability of our method on large-scale real-world corpora. 

\subsection{Adversarial Samples}
We generate adversarial samples in textual form by (1) for each token in the vocabulary, find its closest neighbor after perturbation; (2) for each input sample, replace the token with its found neighbor such that the distance (cosine) between the two is the largest among all tokens and their corresponding neighbors in the input sample. We list some examples on CoLA in Table 5. One may argue that the low evaluation score on CoLA is be expected as the generated adversarial samples are unlikely to be linguistically acceptable. However, these results suggest that the classifier may be confused easily.

\section{Future Work}
In the future, we plan to look into  theoretical explanations for the existence of token-agnostic universal adversarial perturbations. We also seek to find a better way to utilize these universal perturbations, e.g., adversary training. 

\bibliography{emnlp-ijcnlp-2019}
\bibliographystyle{acl_natbib}

\end{document}